# Geolocation Representation from Large Language Models are Generic Enhancers for Spatio-Temporal Learning

### v1

@@ Introduction
The acquisition of universal geolocation representations to improve downstream tasks has been a long-standing pursuit. This paper presents our first attempt to utilize recent advanced LLMs to extract such representations. By the merit of the geospatial knowledge within LLMs, the extracted embedding from the pre-activated layer achieves global coverage and serves as a generic enhancer for spatio-temporal learning. We demonstrate the effectiveness of embedding in various tasks, including GP, LTSF, and GSTF. Empirical results indicate that LLMGeovec can improve the performances of various models simply by incorporating it into the model input (i.e., feature concatenation). In future work, we are interested to see if larger LLMs (e.g., LLaMa3 70B) can further improve the quality of LLMGeovec. We can adopt LLMGeovec in more challenging spatio-temporal learning tasks, such as spatio-temporal imputation (Nie et al. 2024b; Yuan et al. 2022) and traffic flow generation  (Wu et al. 2020a). It is also interesting to explore the possibility of integrating LLMGeovec into pre-trained foundational architectures for unified spatio-temporal learning tasks (Jin et al. 2023; Yuan et al. 2024; Zhang et al. 2024b).
@@ Appendix A Appendices
Descriptions of datasets and models in GTSF

### v2

@@ Introduction
The acquisition of universal geolocation representations to improve downstream tasks has been a long-standing pursuit. This paper presents our first attempt to utilize recent advanced LLMs to extract such representations. By the merit of the geospatial knowledge within LLMs, the extracted embedding from the pre-activated layer achieves global coverage and serves as a generic enhancer for spatio-temporal learning. We demonstrate the effectiveness of embedding in various tasks, including GP, LTSF, and GSTF. Empirical results indicate that LLMGeovec can improve the performances of various models simply by incorporating it into the input (i.e., feature concatenation). In future work, we are interested to see if larger LLMs (e.g., LLaMa3 70B) can further improve the quality of LLMGeovec. We can adopt LLMGeovec in more challenging spatio-temporal learning tasks, such as spatio-temporal imputation (Nie et al. 2024b; Yuan et al. 2022) and traffic flow generation  (Wu et al. 2020a). It is also interesting to explore the possibility of integrating LLMGeovec into pre-trained foundational models for unified spatio-temporal learning (Jin et al. 2023; Yuan et al. 2024; Zhang et al. 2024b).
Acknowledgments
The work described in this paper was supported by grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/15206322 and PolyU/15227424),
@@ Appendix A Appendices
Descriptions of datasets and models in GSTF